OpenStudy (anonymous):

rewrite as a simplified radical (52)^1/2 please help; and explain how to solve ples

OpenStudy (misty1212):

HI

OpenStudy (anonymous):

hello

OpenStudy (misty1212):

the one half in the exponent means to take the square root, that is all

OpenStudy (misty1212):

\[\sqrt{52}\] then i guess you can write that in simplest radical form

OpenStudy (misty1212):

\[52=4\times 13\] \[\sqrt{52}\sqrt{4\times 13}=\sqrt4\sqrt{13}=2\sqrt{13}\]

OpenStudy (anonymous):

i get 2\[\sqrt{13}\]

OpenStudy (misty1212):

most of those steps you do in your head

OpenStudy (misty1212):

yea that's right

OpenStudy (anonymous):

wow ok thank you very much

OpenStudy (misty1212):

you are welcome \(\color\magenta\heartsuit\)\(\color\magenta\heartsuit\)

OpenStudy (anonymous):

u guys can help me with another one?